# Online Smoothed Demand Management

### v1

@@ 2.2 Motivating Applications
Grid-integrated Data Center with Storage. Consider a grid-connected data center with local energy storage. At each time step $t$, the operator decides how much electricity to purchase ($x_{t}$) from the grid at price $p_{t}$, and how much energy to use to meet demand ($z_{t}$); the difference is stored in or drawn from the battery with state $s_{t}$. Base demands $b_{t}$ represent e.g., interactive workloads that must be satisfied immediately, while flexible demands $f_{t}$ capture delay-tolerant batch jobs with deadlines $\Delta_{t}>t$. The operator minimizes their cost of electricity subject to a smoothing penalty ($\mathcal{S}(\cdot)$) that penalizes them for large fluctuations, as sudden shifts in load cause instability and even blackouts on the power grid Kwon:25. Additional penalties model the cost of discharging the battery ($\mathcal{D}(\cdot)$) to account for degradation, promote keeping the battery sufficiently charged to accommodate the risk of an outage, and smooth the rate of internal consumption ($\delta|z_{t}-z_{t-1}|$) to reduce wear-and-tear on components. Unlike prior demand-management formulations Yang:20, OSDM jointly captures both flexible and inflexible workloads and the need for smooth grid- and demand-side decisions. This application is the focus of our case study in § 7.
@@
Thermal Energy Demand Management. An operator manages a thermal energy system (for example, district heating) with a local storage tank NREL:ThermalStorage. At each time step $t$, the operator decides how much thermal energy to purchase ($x_{t}$) at price $p_{t}$ and how much to deliver to meet demand ($z_{t}$); the difference is stored in or drawn from the tank with state $s_{t}$. Base demands $b_{t}$ represent immediate thermal needs, while flexible demands $f_{t}$ can be deferred until a deadline $\Delta_{t}>t$. The operator minimizes the total cost to procure and deliver energy $p_{t}x_{t}+\mathcal{D}(\cdot)$, subject to a smoothing penalty that models power grid stability needs (e.g., if electricity is used to generate thermal energy), since demand spikes can cause grid failure FT:25:europe. An optional penalty on the delivery side $\delta|z_{t}-z_{t-1}|$ models wear-and-tear on thermal components.
@@
Just-in-Time Manufacturing with Material Inventory. A factory converts raw materials into finished products using on-site inventory. At each step $t$, it decides how much material to purchase ($x_{t}$) at price $p_{t}$ and how much product to produce/deliver ($z_{t}$), maintaining warehouse state $s_{t}$. Base orders $b_{t}$ must be filled immediately, while flexible orders $f_{t}$ can be deferred until deadline $\Delta_{t}>t$. The objective is to minimize the purchase and processing cost $p_{t}x_{t}+\mathcal{D}(\cdot)$, while smoothing the purchasing rate ($\mathcal{S}(\cdot)$) and production rate ($\delta|z_{t}-z_{t-1}|$).
@@
Flow Battery Storage Management. A large electricity consumer operates a grid-connected flow battery to shift energy usage Breeze:19. At each time $t$, the operator purchases electricity ($x_{t}$) at price $p_{t}$ and delivers electricity to meet demand ($z_{t}$), with battery charge state $s_{t}$. The goal is to minimize the total cost $p_{t}x_{t}$ subject to a smoothing penalty informed by grid considerations as above, and accounting for delivery inefficiency through $\mathcal{D}(\cdot)$, which captures system losses and auxiliary energy needs.
2.3 Modeling Smoothness Penalty
@@
In OSDM, the smoothness penalty $\mathcal{S}(\cdot)$ in Equation 1 captures an objective to make “smooth” purchasing decisions over time. This is often an important goal (see, e.g., the applications in § 2.2); for instance, industrial consumers like data centers can have electricity consumption patterns with large fluctuations that destabilize the grid and reduce power quality Bloomberg:24. Motivated by these considerations, we consider two natural models for the smoothness penalty $\mathcal{S}(\cdot)$. First, we consider a switching cost model (with corresponding problem referred to as OSDM-S, hereinafter), where the penalty charges the operator proportionally to the change in purchasing decisions:
@@
Second, we consider a tracking cost model (OSDM-T, hereinafter), where the goal is not just to smooth the purchasing rate, but to follow an externally-provided target signal $\{a_{t}\}_{t\in[T]}$. This is motivated by recent interest in programs where grid operators provide load shaping signals to large electricity consumers to encourage consumption patterns that aid grid stability Jiang:14:LoadShaping; Aryandoust:17:LoadShaping. This target signal may be known in advance (e.g., day-ahead) or revealed online (e.g., real-time), and the tracking cost penalizes the operator based on the deviation from this target signal.
@@
2.4 Modeling Delivery Cost
The delivery cost $\mathcal{D}(\cdot)$ in Equation 1 represents the additional expense incurred when delivering energy to the consumer. While often negligible ($\mathcal{D}(\cdot)=0$), it can vary significantly across applications. In a thermal energy system, for instance, efficiency improves when the reservoir (i.e., storage) is fuller or when the temperature differential is small, implying a delivery cost that decreases with the storage state and depends on energy price due to pumping or heat-exchange losses Ruhnau:19. In a just-in-time manufacturing context, $\mathcal{D}(\cdot)$ may reflect a fixed processing cost independent of storage or price. On the other hand, in a flow battery storage system, using stored energy requires some additional input energy to drive pumps and circulate liquid Breeze:19, yielding a cost that increases with discharge volume and depends on the contemporaneous energy price. Given this diversity of cases, we model a general class of monotone, price-dependent delivery costs as follows.
@@ 2.4 Modeling Delivery Cost
Let $c\geq 0$ and $\varepsilon\geq 0$ be delivery cost coefficients known a priori, where $c+\varepsilon\leq 1$. We say that the delivery cost $\mathcal{D}(\cdot)$ in (1) is a monotone affine price-dependent delivery cost if it satisfies one of the following:
@@
    (Increasing case) Let $\mathcal{D}(z_{t},s_{t-1},p_{t})\coloneqq(c\nicefrac{{s_{t-1}}}{{S}}+\varepsilon)p_{t}z_{t}$ – this delivery cost increases linearly from $\varepsilon p_{t}z_{t}$ to $(c+\varepsilon)p_{t}z_{t}$ as the storage state increases from $s_{t-1}=0$ to $s_{t-1}=S$.
@@
    (Decreasing case) Let $\mathcal{D}(z_{t},s_{t-1},p_{t})\coloneqq(c-c\nicefrac{{s_{t-1}}}{{S}}+\varepsilon)p_{t}z_{t}$ – this delivery cost decreases linearly from $(c+\varepsilon)p_{t}z_{t}$ to $\varepsilon p_{t}z_{t}$ as the storage state increases from $s_{t-1}=0$ to $s_{t-1}=S$.
Note that a constant price-dependent delivery cost is captured as a special case when $c=0$.
In the rest of this paper, we present results for the OSDM-S and OSDM-T problems using the decreasing monotone delivery cost case defined in Def. 2.4(b), because it is the most challenging case to analyze: in Theorem 5.2, we show that given the same values of $c$ and $\varepsilon$, the best achievable competitive ratios for the increasing case in Def. 2.4(a) imply it is a strictly easier problem than the decreasing case. Thus, our results primarily focus on this “hard case” and hold for the general monotone definition in Def. 2.4. We note that setting the delivery cost according to Def. 2.4(b) with $c>0$ results in a non-convex (bilinear) objective in (1) due to the dependence on $s_{t-1}$. [^2] The online problem is unaffected by non-convexity because the state $s_{t-1}$ is a fixed value at time $t$ (not a decision variable).
2.5 Assumptions and Notation
In this section, we summarize our main assumptions and additional notation for OSDM. First, we assume that the market prices are bounded: $p_{t}\in[p_{\min},p_{\max}]\ \forall t\in[T]$, where $0<p_{\min}\leq p_{\max}$. $p_{\min}$ and $p_{\max}$ are known to the operator a priori. This is a standard assumption in the literature on online search and related problems such as online knapsack ElYaniv:01; Zhou:08; Larsen:10. It is known that without such an assumption, no online algorithm can achieve a bounded competitive ratio ElYaniv:01; Lorenz:08.
We also assume that it is always feasible to satisfy demand as it becomes due, i.e., if necessary, it is always possible to set $x_{t}=b_{t}+\sum_{\tau:\Delta_{\tau}=t}f_{\tau}$ and $z_{t}=b_{t}+\sum_{\tau:\Delta_{\tau}=t}f_{\tau}$ for any $t\in[T]$. This is a natural assumption that ensures the problem is well-posed – without such an assumption, an online algorithm can be backed into a corner where feasibility becomes impossible. While we do not consider rate constraints in our setting of OSDM (i.e., constraints on the amount of energy that can be purchased or delivered at any time), this gives an implicit condition relevant to such constraints–see discussion of future work in § 8 for more on this point.
@@ 2.5 Assumptions and Notation
As given in Def. 2.4, we assume the delivery cost coefficients $c\geq 0$ and $\varepsilon\geq 0$ satisfy $0\leq c+\varepsilon\leq 1$. This ensures that the delivery cost is not “too large” relative to the purchasing cost, and is realistic in the target applications of OSDM, where the delivery cost (e.g., the added cost to serve demand using stored energy) is a fraction of the original cost to procure said energy.
In OSDM-T, we make two additional assumptions related to the target signal and the tracking cost. First, we assume that the cumulative target is at most the demand, i.e., $\sum_{t=1}^{T}a_{t}\leq\sum_{t=1}^{T}(b_{t}+f_{t})$. This ensures that $\sum_{t\in[T]}a_{t}$ is reasonable relative to the total demand over the horizon. Without this assumption, there are unrealistic OSDM instances with, e.g., a small amount of demand and a consistently large target signal that force any algorithm to pay a large penalty for failing to have enough demand to meet the target. Finally, we assume that the tracking cost coefficient $\eta$ satisfies $\eta\leq\frac{p_{\max}-p_{\min}}{2}$ – this ensures that the tracking cost coefficient is not “too large” relative to the range of market prices. If this assumption is violated (i.e., if $\eta>\frac{p_{\max}-p_{\min}}{2}$), then it is always preferable (from a cumulative purchasing cost perspective) to set $x_{t}=a_{t}$ wherever possible. Thus, we focus on the “interesting regime” where an online algorithm balances the tracking cost and the purchasing cost, allowing for slight deviations from the target signal when they are advantageous.
@@
3 Warmup and Preliminary Results
In this section, we start by discussing similarities between OSDM and related problems in the literature. We review one such problem and an existing algorithm for it, showing that an extension of this algorithm to OSDM fails to achieve a good competitive ratio, even in a simplified setting – this warmup result illustrates the complexity of the problem and motivates observations that we use in the design of our main algorithm, PAAD, described in § 4.
Relation to Online Conversion with Switching Costs. A much-simplified version of the OSDM problem defined in § 2 has been studied under the name of online conversion with switching costs (OCSLechowicz:24. This problem is a special case of OSDM-S where there is a single unit of flexible demand (whose size is known a priori) that must be satisfied by the end of the time horizon (i.e., $f_{0}=1$ and $\Delta_{0}=T$). OCS also does not consider base demand (i.e., $b_{t}=0$ for all $t$), there is no storage (i.e., $S=0$), and there is no delivery cost (i.e., $c=\varepsilon=\delta=0$). The goal is to minimize the total cost of purchasing an asset subject to time-varying prices, while also minimizing a switching cost incurred from changing the purchase amount over time. For the OCS problem, an algorithmic framework called “ramp-on, ramp-off” (RORO) achieves the best possible competitive ratio Lechowicz:24. The key idea underlying RORO is to maintain a single threshold function and completion state (i.e., “how much demand has been satisfied”) that guides the algorithm’s decision-making; this threshold is designed to capture the marginal trade-off between purchasing at the current price and waiting for better prices that may not arrive (see Appendix C for more details on the algorithm and setting).
@@ 3 Warmup and Preliminary Results
We give the full proof of Theorem 3.1 in Appendix D. This result shows that the competitive ratio of the doubling RORO algorithm can be arbitrarily bad – in particular, the first entry in the $\min$ expression above grows to infinity as $\sigma$ approaches one, implying that the competitive ratio of the doubling RORO algorithm is not better than the trivial $O\big(\tfrac{p_{\max}}{p_{\min}}\big)$ bound achievable by any feasible OSDM algorithm. Intuitively, such an extension of RORO cannot account for the demand that arrives in a coarse-grained manner: it fails to resolve the “gap” between how much demand the algorithm anticipates and the cumulative demand that does arrive, and this gap leads to its suboptimality. This limitation motivates the more modular and adaptive design of our main algorithm PAAD, which we present in the next section.
4 A Unified Competitive Algorithm

### v2

@@ 2.2. Motivating Applications
Grid-integrated Data Center with Storage. Consider a grid-connected data center with local energy storage. At each time step $t$, the operator decides how much electricity to purchase ($x_{t}$) from the grid at price $p_{t}$, and how much energy to use to meet demand ($z_{t}$); the difference is stored in or drawn from the battery with state $s_{t}$. Base demands $b_{t}$ represent e.g., interactive workloads that must be satisfied immediately, while flexible demands $f_{t}$ capture delay-tolerant batch jobs with deadlines $\Delta_{t}>t$. The operator minimizes their cost of electricity subject to a smoothing penalty ($\mathcal{S}(\cdot)$) that penalizes them for large fluctuations, as sudden shifts in load cause instability and even blackouts on the power grid (Kwon et al., 2025). Additional penalties model the cost of discharging the battery ($\mathcal{D}(\cdot)$) to account for degradation, promote keeping the battery sufficiently charged to accommodate the risk of an outage, and smooth the rate of internal consumption ($\delta|z_{t}-z_{t-1}|$) to reduce wear-and-tear on components. Unlike prior demand-management formulations (Yang et al., 2020), OSDM jointly captures both flexible and inflexible workloads and the need for smooth grid- and demand-side decisions. This application is the focus of our case study in § 7.
@@
Thermal Energy Demand Management. An operator manages a thermal energy system (for example, district heating) with a local storage tank (National Renewable Energy Laboratory, 2024). At each time step $t$, the operator decides how much thermal energy to purchase ($x_{t}$) at price $p_{t}$ and how much to deliver to meet demand ($z_{t}$); the difference is stored in or drawn from the tank with state $s_{t}$. Base demands $b_{t}$ represent immediate thermal needs, while flexible demands $f_{t}$ can be deferred until a deadline $\Delta_{t}>t$. The operator minimizes the total cost to procure energy $p_{t}x_{t}$ and deliver energy (e.g., using heat exchangers) $\mathcal{D}(\cdot)$, all subject to a smoothing penalty that models power grid stability needs (e.g., if electricity is used to generate thermal energy), since demand spikes can cause grid failure (Mooney et al., 2025). An optional penalty on the delivery side $\delta|z_{t}-z_{t-1}|$ models wear-and-tear on thermal components.
@@
Just-in-Time Manufacturing with Material Inventory. A factory converts raw materials into finished products using on-site inventory. At each step $t$, it decides how much material to purchase ($x_{t}$) at price $p_{t}$ and how much product to produce/deliver ($z_{t}$), maintaining warehouse state $s_{t}$. Base orders $b_{t}$ must be filled immediately, while flexible orders $f_{t}$ can be deferred until deadline $\Delta_{t}>t$. The objective is to minimize a purchase cost and a fixed per-unit processing cost $p_{t}x_{t}+\mathcal{D}(\cdot)$, while smoothing the purchasing rate ($\mathcal{S}(\cdot)$) and production rate ($\delta|z_{t}-z_{t-1}|$).
@@
Flow Battery Storage Management. A large electricity consumer operates a grid-connected flow battery to shift energy usage (Breeze, 2019). At each time $t$, the operator purchases electricity ($x_{t}$) at price $p_{t}$ and delivers electricity to meet demand ($z_{t}$), with battery charge state $s_{t}$. The goal is to minimize the total cost $p_{t}x_{t}$ subject to a smoothing penalty informed by grid considerations as above, and accounting for system losses and auxiliary energy needs through $\mathcal{D}(\cdot)$.
2.3. Modeling Smoothness Penalty
@@
In OSDM, the smoothness penalty $\mathcal{S}(\cdot)$ in Equation 1 captures an objective to make “smooth” purchasing decisions over time. This is often an important goal (see, e.g., the applications in § 2.2); for instance, industrial consumers like data centers can have electricity consumption patterns with large fluctuations that destabilize the grid and reduce power quality (Nicoletti et al., 2024). Motivated by these considerations, we consider two natural models for the smoothness penalty $\mathcal{S}(\cdot)$. First, we consider a switching cost model (with corresponding problem referred to as OSDM-S, hereinafter), where the penalty charges the operator proportionally to the change in purchasing decisions:
@@
Second, we consider a tracking cost model (OSDM-T, hereinafter), where the goal is not just to smooth the purchasing rate, but to follow an externally-provided target signal $\{a_{t}\}_{t\in[T]}$. This is motivated by recent interest in programs where grid operators provide load shaping signals to large electricity consumers to encourage consumption patterns that aid grid stability (Jiang et al., 2014; Aryandoust and Lilliestam, 2017). This target signal may be known in advance (e.g., day-ahead) or revealed online (e.g., real-time), and the tracking cost penalizes the operator based on the deviation from this target signal.
@@
We remark that the OSDM objective in Equation 1 models the smoothness penalty for the delivery cost (i.e., decisions $\{z_{t}\}_{t\in[T]}$) as a fixed switching cost (with coefficient $\delta$). In principle, this penalty could also be modeled as a tracking cost, and we believe that analogous results are attainable in such a model. We focus on the case where specifically the purchasing decisions are subject to a tracking cost for brevity—in a motivating application such as electricity purchasing, the grid operator provides suitable signals (e.g., reserve capacity signals (Colangelo et al., 2025)) to realize this model (on the purchasing side) in practice.
2.4. Modeling Delivery Cost
The delivery cost $\mathcal{D}(\cdot)$ in Equation 1 represents the additional expense incurred when delivering energy to the consumer. While often negligible ($\mathcal{D}(\cdot)=0$), it can vary significantly across applications. In a thermal energy system, for instance, efficiency improves when the reservoir (i.e., storage) is fuller or when the temperature differential is small, implying a delivery cost that decreases with the storage state and depends on energy price due to pumping or heat-exchange losses (Ruhnau et al., 2019). In a just-in-time manufacturing context, $\mathcal{D}(\cdot)$ may reflect a fixed processing cost independent of storage or price. On the other hand, in a flow battery storage system, using stored energy requires some additional input energy to drive pumps and circulate liquid (Breeze, 2019), yielding a cost that increases with discharge volume and depends on the contemporaneous energy price.
To capture this diversity of cases across multiple applications, in what follows, we define a class of monotone and price-dependent delivery costs. We believe this form of delivery cost to be a particularly challenging case that, if analyzed, offers corollary insights for other delivery cost models. It is defined as follows:
@@ 2.4. Modeling Delivery Cost
Let $c\geq 0$ and $\varepsilon\geq 0$ be delivery cost coefficients known a priori. We say that the delivery cost $\mathcal{D}(\cdot)$ in (1) is a monotone affine price-dependent delivery cost if it satisfies one of the following:
@@
    (Increasing case) Let $\mathcal{D}(z_{t},s_{t-1},p_{t})\coloneqq(\frac{cs_{t-1}}{S}+\varepsilon)p_{t}z_{t}$— this delivery cost increases linearly from $\varepsilon p_{t}z_{t}$ to $(c+\varepsilon)p_{t}z_{t}$ as the storage state increases from $s_{t-1}=0$ to $s_{t-1}=S$.
@@
    (Decreasing case) Let $\mathcal{D}(z_{t},s_{t-1},p_{t})\coloneqq(c-\frac{cs_{t-1}}{S}+\varepsilon)p_{t}z_{t}$— this delivery cost decreases linearly from $(c+\varepsilon)p_{t}z_{t}$ to $\varepsilon p_{t}z_{t}$ as the storage state increases from $s_{t-1}=0$ to $s_{t-1}=S$.
Several alternative delivery cost models are “covered by” Def. 2.4—for instance, setting $c=0$ recovers a price-dependent delivery cost that is constant as a function of the storage state. Similarly, a reasonable alternative model might be a delivery cost that is state-dependent but doesn’t depend on the price $p_{t}$—such a model is captured as a simpler sub-case of Def. 2.4, and our algorithm (see § 4) would directly capture any such simpler cases. In the rest of the paper, we primarily give results for the OSDM-S and OSDM-T problems using the decreasing monotone delivery cost case defined in Def. 2.4(b), because it is the most challenging case to analyze. Theorem 5.3 shows that given the same values of $c$ and $\varepsilon$, the best achievable competitive ratios for the increasing case in Def. 2.4(a) imply it is a strictly easier problem than the decreasing case. Thus, our results primarily focus on this “hard case,” holding for the general monotone definition in Def. 2.4 and any simpler subcases of it. We note that setting the delivery cost according to Def. 2.4(b) with $c>0$ results in a non-convex (bilinear) objective in (1) due to the dependence on $s_{t-1}$. [^2] The online problem is unaffected by non-convexity because the state $s_{t-1}$ is a fixed value at time $t$ (not a decision variable).
2.5. Assumptions and Notation
In this section, we summarize our main assumptions and additional notation for OSDM. First, we assume that the market prices are bounded: $p_{t}\in[p_{\min},p_{\max}]\ \forall t\in[T]$, where $0<p_{\min}\leq p_{\max}$. $p_{\min}$ and $p_{\max}$ are known to the operator a priori. This is a standard assumption in the literature on online search and related problems such as online knapsack (El-Yaniv et al., 2001; Zhou et al., 2008; Larsen and Wøhlk, 2010). It is known that without such an assumption, no online algorithm can achieve a bounded competitive ratio (El-Yaniv et al., 2001; Lorenz et al., 2008).
We also assume that it is always feasible to satisfy demand as it becomes due, i.e., if necessary, it is always possible to set $x_{t}=b_{t}+\sum_{\tau:\Delta_{\tau}=t}f_{\tau}$ and $z_{t}=b_{t}+\sum_{\tau:\Delta_{\tau}=t}f_{\tau}$ for any $t\in[T]$. This is a natural assumption that ensures the problem is well-posed—without such an assumption, an online algorithm can be backed into a corner where feasibility becomes impossible. While we do not consider rate constraints in our setting of OSDM (i.e., constraints on the amount of energy that can be purchased or delivered at any time), this gives an implicit condition relevant to such constraints—see discussion of future work in § 8 for more on this point.
@@ 2.5. Assumptions and Notation
Using the notation of Def. 2.4, we assume the delivery cost coefficients $c\geq 0$ and $\varepsilon\geq 0$ satisfy $0\leq c+\varepsilon\leq 1$. This ensures that the delivery cost is not “too large” relative to the purchasing cost, and is realistic in the target applications of OSDM, where the delivery cost (e.g., the added cost to serve demand using stored energy) is a fraction of the original cost to procure said energy.
In OSDM-T, we make two additional assumptions related to the target signal and the tracking cost. First, we assume that the cumulative target is at most the demand, i.e., $\sum_{t=1}^{T}a_{t}\leq\sum_{t=1}^{T}(b_{t}+f_{t})$. This ensures that $\sum_{t\in[T]}a_{t}$ is reasonable relative to the total demand over the horizon. Without this assumption, there are unrealistic OSDM instances with, e.g., a small amount of demand and a consistently large target signal that force any algorithm to pay a large penalty for failing to have enough demand to meet the target. Finally, we assume that the tracking cost coefficient $\eta$ satisfies $\eta\leq\frac{p_{\max}-p_{\min}}{2}$—this ensures that the tracking cost coefficient is not “too large” relative to the range of market prices. If this assumption is violated (i.e., if $\eta>\frac{p_{\max}-p_{\min}}{2}$), then it is always preferable (from a cumulative purchasing cost perspective) to set $x_{t}=a_{t}$ wherever possible. Thus, we focus on the “interesting regime” where an online algorithm balances the tracking cost and the purchasing cost, allowing for slight deviations from the target signal when they are advantageous.
@@
3. Warmup and Preliminary Results
In this section, we start by discussing similarities between OSDM and related problems in the literature. We review one such problem and an existing algorithm for it, showing that an extension of this algorithm to OSDM fails to achieve a good competitive ratio, even in a simplified setting—this warmup result illustrates the complexity of the problem and motivates observations that we use in the design of our main algorithm, PAAD, described in § 4.
Relation to Online Conversion with Switching Costs. A much-simplified version of the OSDM problem defined in § 2 has been studied under the name of online conversion with switching costs (OCS(Lechowicz et al., 2024). This problem is a special case of OSDM-S where there is a single unit of flexible demand (whose size is known a priori) that must be satisfied by the end of the time horizon (i.e., $f_{0}=1$ and $\Delta_{0}=T$). OCS also does not consider base demand (i.e., $b_{t}=0$ for all $t$), there is no storage (i.e., $S=0$), and there is no delivery cost (i.e., $c=\varepsilon=\delta=0$). The goal is to minimize the total cost of purchasing an asset subject to time-varying prices, while also minimizing a switching cost incurred from changing the purchase amount over time. For the OCS problem, an algorithmic framework called “ramp-on, ramp-off” (RORO) achieves the best possible competitive ratio (Lechowicz et al., 2024). The key idea underlying RORO is to maintain a single threshold function and completion state (i.e., “how much demand has been satisfied”) that guides the algorithm’s decision-making; this threshold is designed to capture the marginal trade-off between purchasing at the current price and waiting for better prices that may not arrive (see Appendix C for more details on the algorithm and setting).
@@ 3. Warmup and Preliminary Results
We give the full proof of Theorem 3.1 in Appendix D. This result shows that the competitive ratio of the doubling RORO algorithm can be arbitrarily bad—in particular, the first entry in the $\min$ expression above grows to infinity as $\sigma$ approaches one, implying that the competitive ratio of the doubling RORO algorithm is not better than the trivial $O\big(\tfrac{p_{\max}}{p_{\min}}\big)$ bound achievable by any feasible OSDM algorithm. Intuitively, a simple extension of RORO cannot account for the demand that arrives in a coarse-grained manner: it fails to resolve the “gap” between the demand the algorithm anticipates and the actual demand that arrives, and this gap leads to its suboptimality. This limitation motivates the modular and adaptive design of our main algorithm PAAD, which we present next.
4. A Unified Competitive Algorithm